OpenStudy (anonymous):

Which number is the solution of m + 90 = 150? 50 60 140 240

undeadknight26 (undeadknight26):

easy if solving for m it would be m + 90 = 150 -90 -90 m = 60

OpenStudy (anonymous):

yep! the undeadknight26 is correct

undeadknight26 (undeadknight26):

I like the "the" :D thank you :D